# PDE Generalization of In-Context Operator Networks: A Study on 1D Scalar Nonlinear Conservation Laws

### 14 Jan 2024 v1

@@ Abstract
Can we build a single large model for a wide range of PDE-related scientific learning tasks? Can this model generalize to new PDEs, even of new forms, without any fine-tuning? In-context operator learning and the corresponding model In-context Operator Networks (ICON) [1] represent an initial exploration of these questions. The capability of ICON regarding the first question has been demonstrated in [1]. In this paper, we explore the second question by investigating the generalization capabilities of ICON for conservation laws, a family of PDEs with temporal evolution. We show the positive answer to the second question, i.e., ICON can generalize well to some PDEs with new forms without any fine-tuning. We also show how to broaden the range of problems that ICON can address, by transforming functions and equations to ICON’s capability scope. We believe that the progress in this paper is a significant step towards the goal of training a foundation model for PDE-related tasks under the in-context operator learning framework.
@@ 1 Introduction
This paper aims to further explore the generalization capabilities of ICON for PDEs. In particular, we focus on conservation laws, a family of PDEs with temporal evolution. Our investigation encompasses both time forward and reverse operators. The forward operator takes the initial state as the condition and predicts the system’s future state as the QoI, while the reverse operator is conceptualized by making time-reverse predictions, i.e., inverting the roles of the condition and the QoI. Clearly, varying PDEs, geometries, boundary conditions, and temporal intervals define distinct operators.
@@ 2.2 Forward and Reverse Operators
Each condition or QoI function $u(t,\cdot)$ is represented by a set of key-value pairs, each pair representing a token in the input sequence of the ICON model. In this paper, the key is the spatial coordinate $x$, and the value is the average value of the solution $u$ over the interval $[x,x+\Delta x]$ with $\Delta x=0.01$, for $x=0,0.01,...,0.99$, i.e. 100 key-value pairs or tokens in total.
@@ 2.4 Inference
In contrast, for video prediction, the model is only able to make predictions for $\Delta t$ time ahead with one forward pass. Therefore, one needs to call the model $n/2$ times to make the same prediction, which not only incurs more computational cost, but also makes the model more prone to error accumulation. Even if the model is trained with flexible strides following the video prediction approach, then in order to predict $u((n+n/2)\Delta t)$ with one forward pass, the model would take $(u(0),u(n/2\Delta t),u(n\Delta t))$ as the input sequence, which only has three frames that may not be sufficient to make accurate predictions.
@@ 4.2 In-Distribution Operators
Figure 4: Averaged error v.s. the number of examples used for in-context operator learning.
@@
In Figure 4, we show the forward error and reverse error averaged over the operators on the grid and 100 instances of in-context operator learning for each operator. For all three configurations, it’s clear that both errors decay as the number of in-context examples increases. The consistency loss performs the worst in our experiments, although with the largest computational cost. This may be due to that the model cannot serve as an accurate forward operator surrogate before convergence, and thus the consistency loss is not effective. In the future, a more sophisticated training strategy can be developed to improve the performance of consistency loss, e.g., gradually increasing the weight of consistency loss as the model converges. Since this is not the focus of this paper, we leave it as future work.
@@ 4.4 Change of Variables
The errors for forward and reverse predictions are shown in Figure 11. We can see that when $r=1$, i.e. scaling $u$ to $[-1,1]$, the errors are the lowest. Recall the initial condition is sampled from the Gaussian process with mean 0 and variance 1, it is reasonable that $r=1$ works better than “no change of variables”, or $r=2,3$. The errors for $r=0.5$ are higher. We attribute this to that the prediction errors are amplified when transforming back to the original variables.
@@ 4.5 Varying Strides
Figure 12: The error of prediction with different maximum strides. (a,b) the error of prediction for $\partial_{t}u+\partial_{x}(\sin(u)-\cos(u))=0$. (c,d) the error of prediction for $\partial_{t}u+\partial_{x}(3\sin(u)-3\cos(u))=0$. (a,c) the error of forward prediction. (b,d) the error of reverse prediction. For $f=\sin(u)-\cos(u)$, the error increases as the maximum stride decreases from 0.05 to 0.01 due to error accumulation. For $f=3\sin(u)-3\cos(u)$, the error decreases as the maximum stride increases from 0.01 to 0.03 due to error accumulation, and then increases as the maximum stride continues to increase from 0.03 to 0.05 due to the out-of-distribution effect.
@@ 5 Summary
In this paper, we present a detailed demonstration of ICON for PDEs with temporal evolution, using 1D scalar nonlinear conservation laws as an example. By designing the data prompts appropriately, we can use a single ICON model to make forward and reverse predictions for different equations with different strides. We show that an ICON model trained on conservation laws with cubic flux functions can generalize well to some other flux functions of more general forms, without fine-tuning. We also show how to broaden the range of problems that ICON can address, by transforming functions and equations to ICON’s capability scope via change of variables and proper strides. We believe that the progress in this paper is a significant step towards the goal of training a foundation model for scientific machine learning under the in-context operator learning framework.

### 21 Jan 2024 v2

@@ Abstract
Can we build a single large model for a wide range of PDE-related scientific learning tasks? Can this model generalize to new PDEs, even of new forms, without any fine-tuning? In-context operator learning and the corresponding model In-Context Operator Networks (ICON) represent an initial exploration of these questions. The capability of ICON regarding the first question has been demonstrated previously. In this paper, we present a detailed methodology for solving PDE problems with ICON, and show how a single ICON model can make forward and reverse predictions for different equations with different strides, provided with appropriately designed data prompts. We show the positive evidence to the second question, i.e., ICON can generalize well to some PDEs with new forms without any fine-tuning. This is exemplified through a study on 1D scalar nonlinear conservation laws, a family of PDEs with temporal evolution. We also show how to broaden the range of problems that an ICON model can address, by transforming functions and equations to ICON’s capability scope. We believe that the progress in this paper is a significant step towards the goal of training a foundation model for PDE-related tasks under the in-context operator learning framework.
@@ 1 Introduction
This paper aims to (1) present a detailed methodology for PDE forward and reversal predictions under the in-context operator learning framework, and (2) explore the generalization capabilities of ICON for PDEs.
In particular, we focus on conservation laws, a family of PDEs with temporal evolution. Our investigation encompasses both time forward and reverse operators. The forward operator takes the initial state as the condition and predicts the system’s future state as the QoI, while the reverse operator is conceptualized by making time-reverse predictions, i.e., inverting the roles of the condition and the QoI.
@@ 2.2 Forward and Reverse Operators
Each condition or QoI function $u(t,\cdot)$ is represented by a set of key-value pairs, each pair representing a token in the input sequence of the ICON model. In this paper, the key is the spatial coordinate $x$, and the value is the average value of the solution $u$ over the interval $[x,x+\Delta x]$ with $\Delta x=0.01$, for $x=0,0.01,...,0.99$, i.e. 100 key-value pairs or tokens for each function.
@@ 2.4 Inference
To make reverse predictions, given the sequence $(u(-i\Delta t))_{i=0}^{n}$, we can construct condition-QoI examples as
$\langle u(0),u(-\Delta t)\rangle,\langle u(-\Delta t),u(-2\Delta t)\rangle,\cdots,\langle u(-(n-1)\Delta t),u(-n\Delta t)\rangle,$ (13)
and set $u(-n\Delta t)$ as the question condition, so the ICON model will learn the reverse operator and predict $u(-(n+1)\Delta t)$. Similarly for multi-step reverse predictions.
@@ 4.2 In-Distribution Operators
Figure 4: Averaged error v.s. the number of examples used for in-context operator learning. (a) Forward error. (b) Reverse error.
In Figure 4, we show the forward error and reverse error averaged over the operators on the grid and 100 instances of in-context operator learning for each operator. For all three configurations, it’s clear that both errors decay as the number of in-context examples increases. The consistency loss performs the worst in our experiments, although with the largest computational cost. This may be due to that the model cannot serve as an accurate forward operator surrogate before convergence, and thus the consistency loss is not effective. A more sophisticated training strategy can be developed to improve the performance of consistency loss, e.g., gradually increasing the weight of consistency loss as the model converges. Since this is not the focus of this paper, we leave it as future work.
@@ 4.4 Change of Variables
The errors for forward and reverse predictions are shown in Figure 11. We can see that when $r=1$, i.e. scaling $u$ to $[-1,1]$, the errors are the lowest. Recall the initial condition is sampled from the Gaussian process with mean 0 and variance 1, it is reasonable that $r=1$ works better than “no change of variables”, or $r=2,3$. The errors for $r=0.5$ are higher, which can be attributed to that the prediction errors are amplified when transforming back to the original variables.
@@ 4.5 Varying Strides
Figure 12: The error of prediction with different maximum strides. (a,b) the error of prediction for $\partial_{t}u+\partial_{x}(\sin(u)-\cos(u))=0$. (c,d) the error of prediction for $\partial_{t}u+\partial_{x}(3\sin(u)-3\cos(u))=0$. (a,c) the error of forward prediction. (b,d) the error of reverse prediction.
@@ 5 Summary
In this paper, we present a detailed methodology of ICON for PDEs with temporal evolution, using 1D scalar nonlinear conservation laws as an example. By designing the data prompts appropriately, we can use a single ICON model to make forward and reverse predictions for different equations with different strides. We show that an ICON model trained on conservation laws with cubic flux functions can generalize well to some other flux functions of more general forms, without fine-tuning. We also show how to broaden the range of problems that ICON can address, by transforming functions and equations to ICON’s capability scope via change of variables and proper strides. We believe that the progress in this paper is a significant step towards the goal of training a foundation model for scientific machine learning under the in-context operator learning framework.